Suppose that the CAPM is a good model of risk in the stock market. Suppose also that the average excess return on stocks is 10 percent and that the risk-free interest rate is 1 percent. What would you expect to be the return to stocks with each of the following beta coefficients?

a 20.5

b 0.3

c 1.0

d 2.0
Answer #1

CAPM: Expected Return = Risk-Free Rate + Beta x (Excess Return on Stocks) (return above the risk-free rate)

Risk_Free Rate = 1 % and Excess Return = 10 %

(a) Expected Return = 1 + 20.5 x 10 = 206 %

(b) Expected Return = 1 + 0.3 x 10 = 4 % %

(c) Expected Return = 1 + 1 x 10 = 11 %​​​​​​​

(d) Expected Return = 1 + 2 x 10 = 21 %